OpenStudy (anonymous):

Rationalize the denominator of 7 over 2 plus 3 square root of 2

OpenStudy (anonymous):

\[\frac{7}{2+3\sqrt2}\] \[\frac{7}{2+3\sqrt2}\times \frac{2-3\sqrt2}{2-3\sqrt2}\] is a start

OpenStudy (anonymous):

the denominator is \(2^2-(3\sqrt 2)^2=4-18=-14\) the numerator is whatever you get